# Memory-Efficient Acceleration of Block Low-Rank Foundation Models on Resource Constrained GPUs

### v1

@@ 1 Introduction
Large-scale transformer-based foundation models have achieved remarkable success across language understanding, image classification, and generative tasks (vit-iclr2021; llama-arxiv2023; gpt-neurips2020; gpt-arxiv2022; dit-xl). However, their rapid growth in size is increasingly outpacing the capacity of available hardware. For example, Llama-70B (llama-arxiv2023) requires over 140 GB of memory simply to load its weights in half-precision format, yet some of today’s most powerful commercial GPUs provide only 80 GB. Beyond sheer memory constraints, the reliance of transformer models with dense matrix multiplications introduces significant computational and memory bandwidth bottlenecks during inference. These challenges limit the deployment of models at scale and also their accessibility on resource-constrained devices.
@@
A widely adopted strategy to address these bottlenecks is to approximate weight matrices using low-rank factorizations (lowrank-arxiv2021; lowrank-arxiv2023; lowrank-aistats2024). By representing a dense weight matrix as the product of two smaller matrices, the computational and memory complexity of linear layers can be substantially reduced. However, traditional low-rank decompositions often exhibit sharp accuracy degradation at high compression ratios (gblr-iclr2024), which limits their practicality. To overcome this limitation, structured decompositions such as Monarch (monarch-icml22) and BLAST (blast-neurips24) have been proposed. They leverage block low-rank (BLR) structures to capture the underlying representation of weight matrices more effectively, thereby better preserving accuracy while offering memory savings and computational reduction.
@@
Despite these algorithmic advances, the expected end-to-end speedups often fail to materialize in practice on GPUs. Performance on modern GPUs is governed by a roofline model (roofline-2009; roofline-2013) that balances memory bandwidth, peak computational throughput, and arithmetic intensity (i.e., the ratio of operations to memory traffic). Figure 1 illustrates the roofline model of an NVIDIA A40 GPU for 16-bit brain floating-point (BF16) operations. While structured low-rank (LR) decompositions reduce the nominal compute requirements, we first show that they also introduce additional intermediate data movement, particularly in long-sequence scenarios such as the pre-fill stage of large language model (LLM) inference (code_generation_llm-arxiv2024; form_auto_completion_llm-emnlp2023). This shift can move linear layers from the compute-bound regime into the memory-bound regime, creating a gap between algorithmic promise and system-level reality. The issue arises specifically for devices with limited memory subsystems, such as edge GPUs with small L2 caches (4–6 MB) and DRAM based on DDR technology (Jetson Orin Nano) as well as datacenter GPUs (A40). In such cases, (B)LR decompositions paradoxically degrade performance, despite reducing floating-point operations (FLOP) and model size.
@@
In this work, we analyze the performance of LR, Monarch, and BLAST matrix multiplications for efficient transformer-based foundation model inference, with a particular emphasis on long sequences. We identify and characterize key bottlenecks arising from data movement, suboptimal memory layouts, and compiler limitations, and we introduce optimized implementations using Triton (triton-2019), an open-source intermediate language designed for writing efficient GPU kernels. Our proposed kernels exploit partial fusion, operation reordering, and tailored memory layouts to mitigate the overheads of (B)LR structured matrix multiplications in multi-token inference. Through extensive evaluation, we demonstrate that these optimizations deliver substantial performance gains across diverse transformer models, including GPT2-S, Llama-1/7B, and DiT-XL/2 on both server-grade and edge-class GPUs. Our results establish that BLR-based model compression, when paired with hardware-aware optimizations, offers a viable path toward practical deployment of foundation models in resource-constrained environments. Overall, our contributions can be summarized as follows:
@@ 2.1 Weight Structures
Dense weights can be factorized as ${\bm{W}}={\bm{V}}{\bm{U}}$, where ${\bm{V}}\in\mathbb{R}^{i\times r}$ and ${\bm{U}}\in\mathbb{R}^{r\times o}$. Instead of materializing $\bm{W}$}, the factorization is stored and used directly in computation. This reduces parameter count to $r(i+o)$ and computation to $nr(i+o)$ FLOP. The low-rank assumption can cause accuracy degradation if the chosen rank $r$ does not capture the true structure of $\bm{W}$}. In practice, $r\ll i,o$ is selected such that $r(i+o)<io$, yielding both memory and computational savings (lowrank-iclr2025; lowrank-cvpr2020; lowrank-aistats2024).
@@
Introduced by monarch-icml22, Monarch divides a dense weight into $b_{2}\times b_{1}$ blocks, yielding a BLR representation111monarch-icml22 introduces a “transposed” permutation at the output which we omit here for simplicity. with uniform per-block rank $r^{\prime}$. Each block ${\bm{W}}_{l,k}$ is factorized as
@@
Introduced by blast-neurips24, BLAST represents a weight matrix using a generalized BLR structure. Unlike Monarch, each block ${\bm{W}}_{l,k}$ shares a pair of matrices ${\bm{V}}_{l}$ and ${\bm{U}}_{k}$ while retaining a unique diagonal matrix ${\bm{S}}_{l,k}$ such that the block factorization becomes
@@
blast-neurips24 evaluate the impact of replacing dense linear layers in foundation models with low-rank, Monarch, and BLAST layers. Results are reported on language and vision tasks using Llama-7/1B (llama-arxiv2023), GPT2-S (gpt2-2019), ViT-B (vit-iclr2021), and DiT-XL/2 (dit-xl). Language models are evaluated with WikiText-103/2 perplexity and zero-shot classification accuracy on common sense reasoning benchmarks piqa; hellaswag; winogrande; boolq; openbookqa; arc. Vision models are evaluated on ImageNet imagenet classification accuracy. Diffusion models are evaluated by generating images with a DDPM sampler (ddpm) and computing FID, sFID, and IS against 50,000 ImageNet validation images (step size 250) to quantify generation quality. Table 1 summarizes the results where Monarch mostly improves upon low-rank, while BLAST achieves the best accuracy at the same compression factor (CF).
@@ 2.2 GPU Performance
GPUs integrate many CUDA cores for general-purpose parallelism and tensor cores specialized for matrix operations. The memory hierarchy spans high-capacity but high-latency off-chip DRAM and smaller on-chip caches (L1/L2) (jia2018dissecting). On resource-constrained devices such as the Jetson Orin Nano, the L2 cache is only a few MB (jetson), so the large activations of foundation models often spill to DRAM between kernels rather than being reused from cache. Even some data center GPUs like the A40 provide just a 6 MB shared L2 cache (a40). While L2 is not directly programmable, developers can leverage L1 and shared memory to improve locality and hide latency in software (cache).
@@
Computation is done by many parallel threads organized into thread blocks and scheduled across streaming multiprocessors (SMs). Threads access a hierarchical memory system: high-latency global memory (DRAM), lower-latency but non-programmable L2 cache, explicitly managed shared memory (L1/SRAM) per thread block, and per-thread registers. A typical kernel loads data from global memory into registers or shared memory, performs computations, and writes results back. Kernels are usually written in CUDA, offering fine-grained control but requiring detailed knowledge of GPU architecture to achieve high performance (che2008performance). For example, uncoalesced accesses reduce bandwidth efficiency, making coalesced loads a key optimization (ryoo2008optimization). While NVIDIA’s proprietary libraries often outperform custom CUDA kernels, recent alternatives like OpenAI’s Triton (triton-2019) allow developers to write efficient GPU kernels in a Python-like syntax, with the compiler handling optimizations such as shared memory management, warp scheduling, and memory coalescing (zhou2025linear; li2025tritonbenchbenchmarkinglargelanguage).
@@ 3.1 Single-Token Inference
During single-token inference, typical in the decoding stage of LLMs where $n=1$, the problem becomes memory-bound (llm_inference-2024). In this setting, memory traffic is dominated by weight movement rather than activations. As a result, compressing weights (e.g., by $2\times$) can nearly double throughput. This trend is evident in Llama-7B layers on the A40 GPU (Figure 3, left). Here, (B)LR methods such as BLAST, Monarch, and traditional low-rank all achieve similar performance compared to dense since the bottleneck lies in weight data movement rather than compute.
@@ 4 Memory-Efficient Kernels
Kernel fusion integrates consecutive kernels into one. For Monarch and BLAST, we first explored fully fusing the $\mathtt{bmm}$ kernels, building on prior work for low-rank layers (tensor_fusion-access2024; tensor_fusion-swe2025). In Triton, matrix multiplications are parallelized via 2-D output tiling, where each thread block iterates over the inner dimension and loads operand tiles into shared memory for tensor core computation via the $\mathtt{dot()}$ operator. In the context of full fusion, this tiling leads to redundant weight loads and recomputation of intermediates, often making fusion slower than launching separate kernels. Using 1-D tiles avoids redundancy but restricts rank and parallelism, yielding speedups only for very small ranks (e.g., $\leq 128$) that correspond to extreme compression ratios (e.g., $\geq 8\times$ for $\mathtt{Q/K/V/O_{proj}}$ in Llama-7B). An evaluation of traditional low-rank is provided in Appendix A.1, which shows that larger ranks are slower than dense or fail due to shared memory limits, while small-rank gains quickly diminish with output size due to limited parallelism. Given these limitations, we turn to partial fusion, where only permutations or subsets of $\mathtt{bmm}$ are fused to reduce memory traffic. We next outline kernel-specific optimizations for Monarch and BLAST, whose different parameterizations motivate distinct strategies.
@@ 14 Z’’[:, n * $t_{n}$ : (n + 1) * $t_{n}$, r * $t_{r}$ : (r + 1) * $t_{r}$] = z’’ ()
Permutation-only fusion with tensor core optimization. The earlier strategy (❹) mapped the second $\mathtt{bmm}$ to CUDA cores rather than tensor cores, sacrificing up to $16\times$ higher throughput (flash_attention2-2023). This tradeoff could negate the FLOP savings from BLAST. To preserve tensor-core execution, one alternative is to eliminate only the costly permutations, but this is challenging because BLAST swaps the outer and innermost dimensions. Directly writing into the target layout of the next $\mathtt{bmm}$ leads to uncoalesced stores, as the batch dimension split across thread blocks cannot share data. Our key insight is that transposing the $\mathtt{dot()}$ output before storing is inexpensive in Triton. We therefore reorder the computation as shown in Figure 7. Instead of right-multiplying by $\mathcal{S}$ and $\mathcal{U}$, we transpose their first and last dimensions to obtain $\mathcal{S}^{T}$ and $\mathcal{U}^{T}$, multiply from the left, and transpose intermediate output tiles within each kernel. This keeps $n$ contiguous, while $r$, $b_{1}$, and $b_{2}$ are successively exposed as the batch dimensions across three kernels, each implementing a transposed $\mathtt{bmm}$ with outer-dimension reordering. Reordering eliminates permutation overhead and maintains high tensor-core utilization via Triton’s $\mathtt{dot()}$, a level of efficiency that, to our knowledge, neither $\mathtt{einsum}$ nor PyTorch compiler-guided methods can match.
@@ 5 Experiments and Results
We evaluate our memory-efficient kernels against baseline implementations from the BLAST (blast-neurips24) and Monarch (monarch-icml22) repositories. Prior work focused on accuracy across a variety of language and vision domain tasks. As shown in Table 1, low-rank performs worst, Monarch improves upon low-rank, and BLAST achieves the highest accuracy with the same model compression factor (CF, ranging from $1.8$ to $3\times$). Our evaluation complements these results with detailed performance benchmarking on the same set of models while additionally including Llama-3.2-1B for broader coverage. Details regarding Llama-3.2-1B training are provided in Appendix A.3 with accuracies reported in Table 1.
@@
Nevertheless, BLAST-based models using ❺ achieve substantial acceleration over models that use its baseline and the Monarch baseline. Relative to the BLAST baseline, ❺ provides $3.05\times$ speedup on Llama-7B/A40, $2.48\times$ on Llama-3.2-1B/A40, $3.68\times$ on Llama-3.2-1B/Jetson, and $1.36\times$ on GPT2-S/Jetson end-to-end. Even in settings where attention dominates like DiT-XL/2 (yuan2024ditfastattn), ❺ reduces overall inference time by $1.4\times$ on the Jetson compared to the BLAST baseline. Some cases such as ViT-B show a different trend where most of the gain comes from applying $\mathtt{torch.compile()}$ to the entire model, which already yields speedups over the dense baseline comparable to those of low-rank. In this setting, ❺ does not provide additional benefit and can even regress performance, while ❶ – ❸ offer little improvement. The configuration itself (with only b=3b{=}3 blocks for BLAST, b=4b{=}4 for Monarch, and rank $128$) is small and particularly well-suited for $\mathtt{torch.compile()}$ to produce an optimized baseline.

### v2

@@ 1 Introduction
Large-scale transformer-based foundation models have achieved remarkable success across language understanding, image classification, and generative tasks (Dosovitskiy et al., 2021; Touvron et al., 2023; Brown et al., 2020; Hoffmann et al., 2022; Peebles and Xie, 2022). However, their rapid growth in size is increasingly outpacing the capacity of available hardware. For example, Llama-70B (Touvron et al., 2023) requires over 140 GB of memory simply to load its weights in half-precision format, yet some of today’s most powerful commercial GPUs provide only 80 GB. Beyond sheer memory constraints, the reliance of transformer models with dense matrix multiplications introduces significant computational and memory bandwidth bottlenecks during inference. These challenges limit the deployment of models at scale and also their accessibility on resource-constrained devices.
@@
A widely adopted strategy to address these bottlenecks is to approximate weight matrices using low-rank factorizations (Huh et al., 2021; Yaras et al., 2023; Kwon et al., 2024). By representing a dense weight matrix as the product of two smaller matrices, the computational and memory complexity of linear layers can be substantially reduced. However, traditional low-rank decompositions often exhibit sharp accuracy degradation at high compression ratios (Lee and Kim, 2024), which limits their practicality. To overcome this limitation, structured decompositions such as Monarch (Dao et al., 2022) and BLAST (Lee et al., 2024) have been proposed. They leverage block low-rank (BLR) structures to capture the underlying representation of weight matrices more effectively, thereby better preserving accuracy while offering memory savings and computational reduction.
@@
Despite these algorithmic advances, the expected end-to-end speedups often fail to materialize in practice on GPUs. Performance on modern GPUs is governed by a roofline model (Williams et al., 2009; Yang et al., 2018) that balances memory bandwidth, peak computational throughput, and arithmetic intensity (i.e., the ratio of operations to memory traffic). Figure 1 illustrates the roofline model of an NVIDIA A40 GPU for 16-bit brain floating-point (BF16) operations. While structured low-rank (LR) decompositions reduce the nominal compute requirements, we first show that they also introduce additional intermediate data movement, particularly in long-sequence scenarios such as the pre-fill stage of large language model (LLM) inference (Jiang et al., 2024; Kaneko and Okazaki, 2023). This shift can move linear layers from the compute-bound regime into the memory-bound regime, creating a gap between algorithmic promise and system-level reality. The issue arises specifically for devices with limited memory subsystems, such as edge GPUs with small L2 caches (4–6 MB) and DRAM based on DDR technology (Jetson Orin Nano) as well as datacenter GPUs (A40). In such cases, (B)LR decompositions paradoxically degrade performance, despite reducing floating-point operations (FLOP) and model size.
@@
In this work, we analyze the performance of LR, Monarch, and BLAST matrix multiplications for efficient transformer-based foundation model inference, with a particular emphasis on long sequences. We identify and characterize key bottlenecks arising from data movement, suboptimal memory layouts, and compiler limitations, and we introduce optimized implementations using Triton (Tillet et al., 2019), an open-source intermediate language designed for writing efficient GPU kernels. Our proposed kernels exploit partial fusion, operation reordering, and tailored memory layouts to mitigate the overheads of (B)LR structured matrix multiplications in multi-token inference. Through extensive evaluation, we demonstrate that these optimizations deliver substantial performance gains across diverse transformer models, including GPT2-S, Llama-1/7B, and DiT-XL/2 on both server-grade and edge-class GPUs. Our results establish that BLR-based model compression, when paired with hardware-aware optimizations, offers a viable path toward practical deployment of foundation models in resource-constrained environments. Overall, our contributions can be summarized as follows:
@@ 2.1 Weight Structures
Dense weights can be factorized as ${\bm{W}}={\bm{V}}{\bm{U}}$, where ${\bm{V}}\in\mathbb{R}^{i\times r}$ and ${\bm{U}}\in\mathbb{R}^{r\times o}$. Instead of materializing $\bm{W}$}, the factorization is stored and used directly in computation. This reduces parameter count to $r(i+o)$ and computation to $nr(i+o)$ FLOP. The low-rank assumption can cause accuracy degradation if the chosen rank $r$ does not capture the true structure of $\bm{W}$}. In practice, $r\ll i,o$ is selected such that $r(i+o)<io$, yielding both memory and computational savings (Wang et al., 2025; Idelbayev and Carreira-Perpinán, 2020; Kwon et al., 2024).
@@
Introduced by Dao et al. (2022), Monarch divides a dense weight into $b_{2}\times b_{1}$ blocks, yielding a BLR representation [^1] Dao et al. (2022) introduces a “transposed” permutation at the output which we omit here for simplicity. with uniform per-block rank $r^{\prime}$. Each block ${\bm{W}}_{l,k}$ is factorized as
@@
Introduced by Lee et al. (2024), BLAST represents a weight matrix using a generalized BLR structure. Unlike Monarch, each block ${\bm{W}}_{l,k}$ shares a pair of matrices ${\bm{V}}_{l}$ and ${\bm{U}}_{k}$ while retaining a unique diagonal matrix ${\bm{S}}_{l,k}$ such that the block factorization becomes
@@
Lee et al. (2024) evaluate the impact of replacing dense linear layers in foundation models with low-rank, Monarch, and BLAST layers. Results are reported on language and vision tasks using Llama-7/1B (Touvron et al., 2023), GPT2-S (Radford et al., 2019), ViT-B (Dosovitskiy et al., 2021), and DiT-XL/2 (Peebles and Xie, 2022). Language models are evaluated with WikiText-103/2 perplexity and zero-shot classification accuracy on common sense reasoning benchmarks Bisk et al. (2020); Zellers et al. (2019); Sakaguchi et al. (2021); Clark et al. (2019); Mihaylov et al. (2018); Clark et al. (2018). Vision models are evaluated on ImageNet Deng et al. (2009) classification accuracy. Diffusion models are evaluated by generating images with a DDPM sampler (Ho et al., 2020) and computing FID, sFID, and IS against 50,000 ImageNet validation images (step size 250) to quantify generation quality. Table 1 summarizes the results where Monarch mostly improves upon low-rank, while BLAST achieves the best accuracy at the same compression factor (CF).
@@ 2.2 GPU Performance
GPUs integrate many CUDA cores for general-purpose parallelism and tensor cores specialized for matrix operations. The memory hierarchy spans high-capacity but high-latency off-chip DRAM and smaller on-chip caches (L1/L2) (Jia et al., 2018). On resource-constrained devices such as the Jetson Orin Nano, the L2 cache is only a few MB (NVIDIA, 2024), so the large activations of foundation models often spill to DRAM between kernels rather than being reused from cache. Even some data center GPUs like the A40 provide just a 6 MB shared L2 cache (NVIDIA, 2022). While L2 is not directly programmable, developers can leverage L1 and shared memory to improve locality and hide latency in software (Choo et al., 2014).
@@
Computation is done by many parallel threads organized into thread blocks and scheduled across streaming multiprocessors (SMs). Threads access a hierarchical memory system: high-latency global memory (DRAM), lower-latency but non-programmable L2 cache, explicitly managed shared memory (L1/SRAM) per thread block, and per-thread registers. A typical kernel loads data from global memory into registers or shared memory, performs computations, and writes results back. Kernels are usually written in CUDA, offering fine-grained control but requiring detailed knowledge of GPU architecture to achieve high performance (Che et al., 2008). For example, uncoalesced accesses reduce bandwidth efficiency, making coalesced loads a key optimization (Ryoo et al., 2008). While NVIDIA’s proprietary libraries often outperform custom CUDA kernels, recent alternatives like OpenAI’s Triton (Tillet et al., 2019) allow developers to write efficient GPU kernels in a Python-like syntax, with the compiler handling optimizations such as shared memory management, warp scheduling, and memory coalescing (Zhou et al., 2025; Li et al., 2025).
@@ 3.1 Single-Token Inference
During single-token inference, typical in the decoding stage of LLMs where $n=1$, the problem becomes memory-bound (Yuan et al., 2024a). In this setting, memory traffic is dominated by weight movement rather than activations. As a result, compressing weights (e.g., by $2\times$) can nearly double throughput. This trend is evident in Llama-7B layers on the A40 GPU (Figure 3, left). Here, (B)LR methods such as BLAST, Monarch, and traditional low-rank all achieve similar performance compared to dense since the bottleneck lies in weight data movement rather than compute.
@@ 4 Memory-Efficient Kernels
Kernel fusion integrates consecutive kernels into one. For Monarch and BLAST, we first explored fully fusing the $\mathtt{bmm}$ kernels, building on prior work for low-rank layers (Sun et al., 2024; Al Awar et al., 2025). In Triton, matrix multiplications are parallelized via 2-D output tiling, where each thread block iterates over the inner dimension and loads operand tiles into shared memory for tensor core computation via the $\mathtt{dot()}$ operator. In the context of full fusion, this tiling leads to redundant weight loads and recomputation of intermediates, often making fusion slower than launching separate kernels. Using 1-D tiles avoids redundancy but restricts rank and parallelism, yielding speedups only for very small ranks (e.g., $\leq 128$) that correspond to extreme compression ratios (e.g., $\geq 8\times$ for $\mathtt{Q/K/V/O_{proj}}$ in Llama-7B). An evaluation of traditional low-rank is provided in Appendix A.1, which shows that larger ranks are slower than dense or fail due to shared memory limits, while small-rank gains quickly diminish with output size due to limited parallelism. Given these limitations, we turn to partial fusion, where only permutations or subsets of $\mathtt{bmm}$ are fused to reduce memory traffic. We next outline kernel-specific optimizations for Monarch and BLAST, whose different parameterizations motivate distinct strategies.
@@ 14 Z’’[:, n * $t_{n}$ : (n + 1) * $t_{n}$, r * $t_{r}$ : (r + 1) * $t_{r}$] = z’’ ()
Permutation-only fusion with tensor core optimization. The earlier strategy (❹) mapped the second $\mathtt{bmm}$ to CUDA cores rather than tensor cores, sacrificing up to $16\times$ higher throughput (Dao, 2023). This tradeoff could negate the FLOP savings from BLAST. To preserve tensor-core execution, one alternative is to eliminate only the costly permutations, but this is challenging because BLAST swaps the outer and innermost dimensions. Directly writing into the target layout of the next $\mathtt{bmm}$ leads to uncoalesced stores, as the batch dimension split across thread blocks cannot share data. Our key insight is that transposing the $\mathtt{dot()}$ output before storing is inexpensive in Triton. We therefore reorder the computation as shown in Figure 7. Instead of right-multiplying by $\mathcal{S}$ and $\mathcal{U}$, we transpose their first and last dimensions to obtain $\mathcal{S}^{T}$ and $\mathcal{U}^{T}$, multiply from the left, and transpose intermediate output tiles within each kernel. This keeps $n$ contiguous, while $r$, $b_{1}$, and $b_{2}$ are successively exposed as the batch dimensions across three kernels, each implementing a transposed $\mathtt{bmm}$ with outer-dimension reordering. Reordering eliminates permutation overhead and maintains high tensor-core utilization via Triton’s $\mathtt{dot()}$, a level of efficiency that, to our knowledge, neither $\mathtt{einsum}$ nor PyTorch compiler-guided methods can match.
@@ 5 Experiments and Results
We evaluate our memory-efficient kernels against baseline implementations from the BLAST (Lee et al., 2024) and Monarch (Dao et al., 2022) repositories. Prior work focused on accuracy across a variety of language and vision domain tasks. As shown in Table 1, low-rank performs worst, Monarch improves upon low-rank, and BLAST achieves the highest accuracy with the same model compression factor (CF, ranging from $1.8$ to $3\times$). Our evaluation complements these results with detailed performance benchmarking on the same set of models while additionally including Llama-3.2-1B for broader coverage. Details regarding Llama-3.2-1B training are provided in Appendix A.3 with accuracies reported in Table 1.
@@
Nevertheless, BLAST-based models using ❺ achieve substantial acceleration over models that use its baseline and the Monarch baseline. Relative to the BLAST baseline, ❺ provides $3.05\times$ speedup on Llama-7B/A40, $2.48\times$ on Llama-3.2-1B/A40, $3.68\times$ on Llama-3.2-1B/Jetson, and $1.36\times$ on GPT2-S/Jetson end-to-end. Even in settings where attention dominates like DiT-XL/2 (Yuan et al., 2024b), ❺ reduces overall inference time by $1.4\times$ on the Jetson compared to the BLAST baseline. Some cases such as ViT-B show a different trend where most of the gain comes from applying $\mathtt{torch.compile()}$ to the entire model, which already yields speedups over the dense baseline comparable to those of low-rank. In this setting, ❺ does not provide additional benefit and can even regress performance, while ❶ – ❸ offer little improvement. The configuration itself (with only b=3b{=}3 blocks for BLAST, b=4b{=}4 for Monarch, and rank $128$) is small and particularly well-suited for $\mathtt{torch.compile()}$ to produce an optimized baseline.